jhonyy9 (jhonyy9):

- there are again same numbers ? 12345679 * 9 ----------------- 111111111 what is your opinion about ? ty.

jhonyy9 (jhonyy9):

@ganeshie8 @Kainui @rebeccaxhawaii @satellite73 @zepdrix @mathstudent55 @TheSmartOne

OpenStudy (mariamikayla):

12345679*9= 111 111 111 9*9=81 7*9=63+8=71 6*9=54+7=61 5*9=45+6=51 4*9=36+5=41 3*9=27+4=31 2*9=18+3=21 1*9=9+2=11